Portobello mushrooms sell for $6.60/kg and shiitake mushrooms sell for $11.00/kg. Find the mass of each type of mushroom in 1-kg bags selling for $8.36.
The answers are 600g portobello and 400g shiitake but I got 300g portobello and 520g shiitake.
Thanks, in advance!!

Answer by amoresroy(333)

Portobello mushrooms sell for $6.60/kg and shiitake mushrooms sell for $11.00/kg. Find the mass of each type of mushroom in 1-kg bags selling for $8.36.
The answers are 600g portobello and 400g shiitake but I got 300g portobello and 520g shiitake.
Let P = mass of Portobello mushrooms in Kg
1-P = mass of Shiitake mushrooms in Kg
Equation:
6.60P + 11.0(1-P) = 8.36
6.6P + 11.0 - 11P = 8.36
-4.4P = 8.36 - 11.0
P = -2.64/-4.4
P = 0.6 = 600 g
1-P = 1 - .6
= 0.4 = 400 gm
So the answers are 600 g Portobello & 400 g Shiitake